# STIQ: Safeguarding Training and Inferencing of Quantum Neural Networks from Untrusted Cloud

### v1

@@ 5.2. Simulation Results
Figure 7. Performance of 8-qubit 2-layer QNN on Fashion-4 dataset. The plot clearly demonstrates that the combined performance of STIQ closely aligns with that of the baseline model, while the individual performances of QNN-1 and QNN-2 show decreased accuracy due to the intentional obfuscation introduced by our approach.
@@
Table 4. Test accuracies for 8-qubit QNNs trained on various datasets and infered across different hardware setups. The baseline, QNN-1, and QNN-2 models are tested on devices replicating the noise characteristics and basis gates of IBM_Sherbrooke, IBM_Hanoi, and IBM_Cairo, respectively.
Penalty Baseline QNN-1 QNN-2 Combined
@@
Letters-4 91.0 37.0 21.0 91.2
Different Datasets: After determining the optimal divergence loss function and penalty value ($\lambda$) for achieving balanced performance, we train larger, more complex 8-qubit and 2-layer (PQC-19 (sim2019expressibility, )) QNNs. For this training, we use CosineSimilarity as our divergence loss function ($L_{D}$), with a penalty $\lambda$ set at 0.3, as identified through our earlier analysis. We conduct extensive training and testing on various datasets to evaluate the effectiveness of our proposed STIQ model under these conditions. The performance metrics, illustrated in Fig. 7, provide a detailed comparison between the various models. The results depicted in the figure demonstrate that the combined performance of our STIQ closely aligns with that of the baseline model, while the individual performances of QNN-1 and QNN-2 show decreased accuracy due to the intentional obfuscation introduced by our STIQ approach. Table 3 shows the test performance of STIQ on different datasets. Eventhough, test loss of the combined model is slightly worse across all the datasets, we can clearly see that the combined performance closely matches to that of the unobfuscated baseline model when comparing the test accuracies. Furthermore, the performance of individual models is also greatly obfuscated demonstrating how effective STIQ is in protecting QNNs from untrusted cloud providers.
Performance under Noise: To evaluate the effectiveness of STIQ in noisy environments, we use our trained 8-qubit 2-layer QNN models on noiseless hardware and then test them on noisy hardware that simulates the noise models and basis gates of real quantum devices. We conducted tests using 300 images under two different setups: (i) all models were executed on a single device that mimics the noise model and basis gates of IBM_Sherbrooke, and (ii) in a scenario that mirrors practical applications more closely, where each model—baseline, QNN-1, and QNN-2—was executed on different devices reflecting the noise characteristics and native gates of IBM_Sherbrooke, IBM_Hanoi, and IBM_Cairo, respectively. This is a more realistic scenario since it helps mitigate the risk of reverse engineering by cloud providers, as it diversifies the hosting environments for the models. Table 4 presents the test results for models trained on various datasets and tested across different devices, showing that STIQ consistently matches the baseline performance with no degradation. Figure 8 illustrates the inference performance of all models, tested either on the a) same or on b) different devices. In both scenarios, the overall performance is comparable to the baseline, with individual models demonstrating significant obfuscation. This highlights STIQ’s effectiveness in maintaining performance integrity under realistic noise conditions.
5.3. Real Hardware Results
Here, we evaluated the performance of STIQ on real quantum hardware by testing 4-qubit 2-layer QNNs (with PQC: circuit-1 (sim2019expressibility, )) on IBM’s 127-qubit IBM_Sherbrooke, as shown in Table 5. The results from the Letters-4 dataset indicate that the combined performance of the models was similar to the baseline, achieving similar accuracy and slightly higher loss. However, the performance of individual QNNs was significantly obfuscated, with an average accuracy reduction of 50% and a loss nearly 8$\times$ that of the baseline and combined predictions. This demonstrates STIQ’s effectiveness in a realistic scenario, where models are inferred on real quantum hardware via a cloud provider. Thus, STIQ not only reduces the risk of adversarial threats but also maintains the integrity and confidentiality of quantum computing models in cloud environments.
5.4. Discussion
Ensemble Learning vs STIQ: Although previous studies have explored the use of ensemble learning in quantum classifiers (schuld2018quantum, ; silver2022quilt, ), the foundational approaches differ significantly from the proposed STIQ methodology. In earlier works, individual classifiers are trained to each make correct predictions independently, with their outputs then aggregated and weighted according to their respective accuracies. In contrast, STIQ employs a novel training procedure where individual classifiers are intentionally trained to make highly inaccurate predictions, yet their combined output results in correct predictions. Moreover, the primary motivation behind prior research was to develop quantum-based ensemble models that deliver robust performance without specifically addressing security concerns. Conversely, STIQ’s methodology is designed fundamentally to protect individual classifiers from cloud-based adversarial attacks while still achieving optimal performance under noise.
@@ 5.4. Discussion
Figure 8. Test performance of 8-qubit QNNs trained on Letters-4 dataset in a noisy environment: (a) Performance across models on a single device with noise and basis gates similar to IBM_Sherbrooke, and (b) Performance across different hardware setups, where the baseline, QNN-1, and QNN-2 are executed on devices mimicking the noise characteristics of IBM_Sherbrooke, IBM_Hanoi, and IBM_Cairo respectively.
@@
Overhead Analysis: Since STIQ involves training atleast 2 QNNs for each task, this increases the total computational overhead by approx 2$\times$ depending on the architecture of the individual QNNs. Using simpler models could help reduce this overhead but this might come at a cost of slightly decreased overall performance. We conducted experiments to evaluate this trade-off by comparing the performance of a baseline 8-qubit, 2-layer QNN (PQC: circuit-8 (sim2019expressibility, )) with that of a smaller 4-qubit, 2-layer QNNs (PQC: circuit-6 (sim2019expressibility, )). Figure 9 presents the train and test accuracies for these models, demonstrating that while smaller models lead to an average performance degradation of about 2.5%, they significantly reduce computational demands. This efficiency allows for training on smaller quantum hardware, such as 5-qubit or 7-qubit systems, which can decrease the overall costs associated with training multiple models required by STIQ. Furthermore, while STIQ does increase the overall computational cost, it does not necessarily lead to a doubling of the training time. This is because the individual QNNs within the STIQ framework can be trained concurrently on different hardware platforms. As a result, if queue depths and gate execution times are similar, the total training time remains comparable to that of a baseline single-model training scenario.
@@
Table 5. Inference performance of 4-qubit 2-layer QNNs on real quantum hardware where all the models are tested on IBM_Sherbrooke.
Inference Baseline QNN-1 QNN-2 Combined
Accuracy 80.0 40.0 20.0 80.0
Loss 0.54 3.57 4.62 0.60
Figure 9. Train and test accuracies for different models on the Kuzushiji-4 dataset, comparing an 8-qubit baseline model with smaller 4-qubit models, QNN-1 and QNN-2.
6. Conclusion
We address the risk of model theft when training or hosting QNNs on untrusted cloud platforms by introducing STIQ that utilizes ensemble techniques. STIQ involves hosting individual QNNs in such a manner that their outputs, when considered separately, are meaningless, effectively rendering them useless to potential attackers. However, when these outputs are aggregated locally, they yield the correct prediction, safeguarding the intellectual property embedded within the QNNs. We rigorously tested our technique across a variety of architectures, employing both noiseless simulators and those mimicking real device noise as well as directly on IBM_Sherbrooke. Our results demonstrate a significant degradation in the accuracy of individual models, while the accuracy of the combined output matches that of an unobfuscated QNN. However, this comes at a cost on increased training and inferencing overhead since it requires users to train multiple models.

### v2

@@ V-A Setup
Figure 7: Performance of 8-qubit 2-layer QNN on Fashion-4 dataset. The plot clearly demonstrates that the combined performance of STIQ closely aligns with that of the baseline model, while the individual performances of QNN-1 and QNN-2 show decreased accuracy due to the intentional obfuscation introduced by our approach.
TABLE IV: Comparison of Obfuscation and Performance Gap ($\Delta$) between STIQ and Baseline QNN Across Different Datasets. Obfuscation is defined using the ratio mean(QNN-1, $\frac{\text{mean(QNN-1, QNN-2)}}{\text{Baseline}}$, indicating the average relative performance of the obfuscated QNNs compared to the Baseline. Performance $\Delta$ is measured as $\frac{\text{Combined}}{\text{Baseline}}$, representing the performance of the combined QNN outputs relative to the Baseline.
Obfuscation Performance $\Delta$
Dataset Accuracy Loss Accuracy Loss
MNIST-4 0.3$\times$ 41.3$\times$ 1.0$\times$ 1.37$\times$
Fashion-4 0.4$\times$ 21.4$\times$ 1.0$\times$ 1.22$\times$
Kuzushiji-4 0.4$\times$ 12.7$\times$ 1.01$\times$ 1.02$\times$
Letters-4 0.3$\times$ 16.5$\times$ 1.01$\times$ 1.15$\times$
V-B Simulation Results
Penalty ($\lambda$) Analysis: To achieve the goal of balancing obfuscation and performance in our QNN setup, we first need to determine the optimal value for the penalty parameter, $\lambda$, as introduced in Eq. 1. This parameter is critical for adjusting the impact of the divergence loss $L_{D}$, in this case, CosineSimilarity, on the overall training process. Our objective is to find a $\lambda$ value that allows individual models to maintain a high degree of obfuscation without significantly compromising the combined performance of the models when compared to an unobfuscated baseline model. An analysis of this is illustrated in Fig. 5 and Table I, which shows the test accuracy of a 4-qubit QNN model trained on the Fashion-4 dataset. The figure reveals that as $\lambda$ increases, enhancing the level of obfuscation by giving more weight to the divergence loss, the performance of the individual models, QNN-1 and QNN-2, begins to decline. This decrease in performance suggests that a high penalty on divergence loss can negatively affect the model’s ability to optimize effectively, as the divergence loss begins to dominate over the classification loss. Based on the trends observed in the figure, it becomes clear that an optimal range for $\lambda$ exists. Specifically, for the CosineSimilarity divergence function, a $\lambda$ value within the range of $[0.1,0.3]$ seems to be the optimal range. Within this range, the models achieve a desirable level of obfuscation (45-75%) while still performing closely to the unobfuscated baseline model in terms of accuracy ($\leq 1\%$).
Divergence Loss ($L_{D}$): Next, we evaluate the effects of employing various divergence loss functions on the training and performance of 4-qubit QNNs on the MNIST-4 dataset. The divergence loss functions under consideration are L1 loss, L2 loss, and CosineSimilarity. The goal is to understand how each loss function influences the model’s test accuracy, as depicted in Fig. 6 and Table II. To ensure effective training of the models with each type of loss function, it was necessary to empirically determine suitable penalty values for each case. This step was crucial because the choice of divergence loss function and its corresponding penalty value directly affects the model’s ability to learn. The analysis revealed a notable pattern: while initially, the models trained with L1 and L2 loss functions showed promising performance, their combined performance began to decline after reaching a certain number of epochs. This decline is is because of the divergence loss ($L_{D}$), which, beyond a certain point, starts to divert more from the overall learning process than the classification loss ($L_{C}$) contributes. Essentially, the models become too focused on minimizing $L_{D}$ at the expense of $L_{C}$, leading to poorer overall performance. In contrast, models utilizing CosineSimilarity as the divergence loss function exhibited a more stable training progression. This stability suggests that CosineSimilarity, unlike L1 or L2 loss, provides a balance that prevents the divergence loss from overwhelming the classification loss. Consequently, our findings indicate that CosineSimilarity is a preferable choice for the divergence loss function in the context of training QNNs, as it supports more consistent and effective learning.
Different Datasets: After determining the optimal divergence loss function and penalty value ($\lambda$) for achieving balanced performance, we train larger, more complex 8-qubit and 2-layer (PQC-19 [31]) QNNs. For this training, we use CosineSimilarity as our divergence loss function ($L_{D}$), with a penalty $\lambda$ set at 0.3, as identified through our earlier analysis. We conduct extensive training and testing on various datasets to evaluate the effectiveness of our proposed STIQ model under these conditions. The performance metrics, illustrated in Fig. 7, provide a detailed comparison between the various models. The results depicted in the figure demonstrate that the combined performance of STIQ closely aligns with that of the baseline model, while the individual performances of QNN-1 and QNN-2 show decreased accuracy due to the intentional obfuscation introduced by our STIQ approach. Table III shows the test performance of STIQ across various datasets. Although the test loss of the combined model is slightly higher across all datasets, the test accuracies indicate that the combined performance closely matches that of the unobfuscated baseline model. This observation is further supported by Table IV, which demonstrates the average obfuscation levels of QNN-1 and QNN-2 relative to the baseline QNN, as well as the performance gap ($\Delta$) between the combined and baseline models. On average, the accuracy and loss of the obfuscated QNNs are 0.35×0.35$\times$0.35 × and 23×23$\times$23 × those of the baseline QNN, respectively. However, the combined accuracy aligns closely (1×1$\times$1 ×) with that of the baseline, albeit with a 1.19×1.19$\times$1.19 × higher loss. These findings further showcases STIQ’s effectiveness in safeguarding QNNs from untrusted cloud providers.
@@ V-B Simulation Results
TABLE V: Test accuracies for 8-qubit QNNs trained on various datasets and infered across different hardware setups. The baseline, QNN-1, and QNN-2 models are tested on devices replicating the noise characteristics and basis gates of IBM_Sherbrooke, IBM_Hanoi, and IBM_Cairo, respectively.
Dataset Baseline QNN-1 QNN-2 Combined
@@
Letters-4 91.0 37.0 21.0 91.0
@@
Figure 8: Test performance of 8-qubit QNNs trained on Letters-4 dataset in a noisy environment: (a) Performance across models on a single device with noise and basis gates similar to IBM_Sherbrooke, and (b) Performance across different hardware setups, where the baseline, QNN-1, and QNN-2 are executed on devices mimicking the noise characteristics of IBM_Sherbrooke, IBM_Hanoi, and IBM_Cairo respectively.
Performance under Noise: To evaluate the effectiveness of STIQ in noisy environments, we use our trained 8-qubit 2-layer QNN models on noiseless hardware and then test them on noisy hardware that simulates the noise models and basis gates of real quantum devices. We conducted tests using 300 images under two different setups: (i) all models were executed on a single device that mimics the noise model and basis gates of IBM_Sherbrooke, and (ii) in a scenario that mirrors practical applications more closely, where each model—baseline, QNN-1, and QNN-2—was executed on different devices reflecting the noise characteristics and native gates of IBM_Sherbrooke, IBM_Hanoi, and IBM_Cairo, respectively. This is a more realistic scenario since it helps mitigate the risk of reverse engineering by cloud providers, as it diversifies the hosting environments for the models. Table V presents the test results for models trained on various datasets and tested across different devices, showing that STIQ consistently matches the baseline performance with no degradation. Figure 8 illustrates the inference performance of all models, tested either on the a) same or on b) different devices. In both scenarios, the overall performance is comparable to the baseline, with individual models demonstrating significant obfuscation. This highlights STIQ’s effectiveness in maintaining performance integrity under realistic noise conditions.
V-C Real Hardware Results
Here, we evaluated the performance of STIQ on real quantum hardware by testing 4-qubit 2-layer QNNs (with PQC: circuit-1 [31]) on IBM_Kyoto, as shown in Table VI. The results from the Letters-4 dataset indicate that the combined performance of the models was similar to the baseline, achieving similar accuracy and slightly higher loss. However, the performance of individual QNNs was significantly obfuscated, with an average accuracy reduction of 60% and a loss nearly 6$\times$ that of the baseline and combined predictions. This demonstrates STIQ’s effectiveness in a realistic scenario, where models are inferred on real quantum hardware via a cloud provider.
TABLE VI: Performance evaluation of QNNs on real quantum hardware. The models were initially tested on IBM_Kyoto. Subsequently they were also tested on different hardwares: Baseline on IBM_Sherbrooke, QNN-1 on IBM_Brisbane, and QNN-2 on IBM_Kyoto.
IBM_Kyoto
Metric Baseline QNN-1 QNN-2 Combined
Accuracy 89.0 40.0 20.0 88.0
Loss 0.53 1.25 3.14 0.61
IBM_Sherbrooke |||| IBM_Brisbane |||| IBM_Kyoto
Metric Baseline QNN-1 QNN-2 Combined
Accuracy 90.0 40.0 20.0 90.0
Loss 0.34 2.77 3.28 0.42
We also considered another practical setting where each model was inferred on distinct quantum hardwares to enhance security by diversifying the hosting hardware for the models. Specifically, the baseline model was tested on IBM_Sherbrooke, QNN-1 on IBM_Brisbane, and QNN-2 on IBM_Kyoto. The results, shown in Table VI, demonstrate that STIQ effectively maintains performance even across different systems. This indicates that STIQ not only mitigates adversarial threats but also preserves the integrity and confidentiality of QML models in cloud environments.
V-D Discussion
Scalability: Proposed STIQ scales well with increase in the size of individual QNNs and dataset. In order to quantify this, we trained QNNs sized 4-qubit to 16-qubit each having 2-layers on a larger MNIST-10 dataset with 2000 samples with train:test of 70:30. We used SPSA for gradient calculation which is a zero-order optimization method known for it’s efficiency and noise-robustness in training QML models when coupled with classical optimizers like Adam [36]. From Table VII, we can clearly see that even with increase in size of QNN, STIQ is effectively able to highly obfuscated individual hosted QNNs. Interestingly, for QNNs trained on a larger dataset i.e. MNIST-10 in this case, STIQ actually outperforms the Baseline QNN by up to 6%, suggesting that STIQ not only protects QNNs from cloud-based adversaries but also leads to superior performance in larger QNNs.
Applicability to VQAs: STIQ can be effectively adapted to enhance the security of other VQAs such as the Quantum Approximate Optimization Algorithm (QAOA) and the Variational Quantum Eigensolver (VQE) against cloud-based adversaries. The loss function ($L$), in problems like finding the MaxCut using QAOA or determining the ground state energy of a molecule using VQE, involves maximizing or minimizing the expectation value of a Hamiltonian ($H$). Consequently, ($L$) in these scenarios can be expressed as $L=\langle H\rangle$. Using this formulation of $L$, the STIQ training methodology can be applied to concurrently optimize two VQA circuits ($U_{1},U_{2}$). Considering each circuit produces obfuscated expectation values, $\langle H_{1}\rangle$ and $\langle H_{2}\rangle$ independently and a divergence function $D()$, penalty $\lambda$ and aggregate function $A()$, the new loss function to optimize both $U_{1}$ and $U_{2}$ would be:
$L=A(\langle H_{1}\rangle\ ,\langle H_{2}\rangle\\ )+\lambda\cdot D(\langle H_{1}\rangle\ ,\langle H_{2}\rangle\\ )$
Thus, STIQ is not merely confined to QNNs; it can be seamlessly extended to any VQA that relies on optimizing parameters based on a defined loss function.
TABLE VII: Test accuracies for different qubit 2-layer QNNs trained on larger MNIST-10 datasets using SPSA for gradient calculation. STIQ seems to outperform Baseline while significantly obfuscating individual QNNs.
Size Baseline QNN-1 QNN-2 Combined
4-Q 57.2 22.3 11.1 65.3
8-Q 69.1 8.33 20.5 75.5
12-Q 89.6 21.5 10.0 93.0
16-Q 85.5 13.8 14.3 90.5
Ensemble Learning vs STIQ: Although previous studies have explored the use of ensemble learning in quantum classifiers [37, 38, 39], the foundational approaches differ significantly from the proposed STIQ methodology. In earlier works, individual classifiers are trained to each make correct predictions independently, with their outputs then aggregated and weighted according to their respective accuracies. In contrast, STIQ employs a novel training procedure where individual classifiers are intentionally trained to make highly inaccurate predictions, yet their combined output results in correct predictions. Moreover, the primary motivation behind prior research was to develop quantum-based ensemble models that deliver robust performance without specifically addressing security concerns. Conversely, STIQ is designed fundamentally to protect individual classifiers from cloud-based adversarial attacks while still achieving optimal performance.
Figure 9: Train and test accuracies for different models on the Kuzushiji-4 dataset, comparing an 8-qubit baseline model with smaller 4-qubit models, QNN-1 and QNN-2.
@@ V-D Discussion
Overhead Analysis: Since STIQ involves training atleast 2 QNNs for each task, this increases the total computational overhead by approx 2$\times$ depending on the architecture of the individual QNNs. Using simpler models could help reduce this overhead but this might come at a cost of slightly decreased overall performance. We conducted experiments to evaluate this trade-off by comparing the performance of a baseline 8-qubit, 2-layer QNN (PQC: circuit-8 [31]) with that of a smaller 4-qubit, 2-layer QNNs (PQC: circuit-6 [31]). Figure 9 presents the train and test accuracies for these models, demonstrating that while smaller models lead to an average performance degradation of about 2.5%, they significantly reduce computational demands. This efficiency allows for training on smaller quantum hardwares, which can decrease the overall costs associated with training multiple models. Furthermore, while STIQ does increase the overall computational cost, it does not necessarily lead to a doubling of the training time. This is because the individual QNNs within the STIQ framework can be trained concurrently on different hardware platforms. As a result, if queue depths and gate execution times are similar, the total training time remains comparable to that of a baseline single-model training scenario.
@@
VI Conclusion
We address the risk of model theft when training or hosting QNNs on untrusted cloud platforms by introducing STIQ that utilizes ensemble techniques. STIQ involves hosting individual QNNs in such a manner that their outputs, when considered separately, are meaningless, effectively rendering them useless to potential attackers. However, when these outputs are aggregated locally, they yield the correct prediction, safeguarding the intellectual property embedded within the QNNs. We rigorously tested our technique across a variety of architectures, employing both noiseless simulators and those mimicking real device noise as well as directly on IBM’s real quantum hardware. Our results demonstrate a significant degradation in the accuracy of individual models, while the accuracy of the combined output matches that of an unobfuscated QNN.